Use De Moivre’s Theorem to determine the cube root of Z and leave your answer in polar


form with the angle in radians


(a) Z = 1+i√3


1
Expert's answer
2022-05-03T16:44:47-0400

a) The polar form of "1+i\sqrt{3}" is

"2\big(\cos(\dfrac{\pi}{3})+i\sin(\dfrac{\pi}{3})\big)"

We have that "r=2, \theta=\dfrac{\pi}{3}, n=3."

According to the De Moivre's Formula

"k=0:"


"\sqrt[3]{2}\big(\cos(\dfrac{\pi/3+2\pi(0)}{3})+i\sin(\dfrac{\pi/3+2\pi(0)}{3})\big)"


"=\sqrt[3]{2}\big(\cos(\dfrac{\pi}{9})+i\sin(\dfrac{\pi}{9})\big)"

"k=1:"


"\sqrt[3]{2}\big(\cos(\dfrac{\pi/3+2\pi(1)}{3})+i\sin(\dfrac{\pi/3+2\pi(1)}{3})\big)"


"=\sqrt[3]{2}\big(\cos(\dfrac{7\pi}{9})+i\sin(\dfrac{7\pi}{9})\big)"

"=\sqrt[3]{2}\big(-\cos(\dfrac{\pi}{9})+i\sin(\dfrac{\pi}{9})\big)"

"k=2:"


"\sqrt[3]{2}\big(\cos(\dfrac{\pi/3+2\pi(2)}{3})+i\sin(\dfrac{\pi/3+2\pi(2)}{3})\big)"


"=\sqrt[3]{2}\big(\cos(\dfrac{13\pi}{9})+i\sin(\dfrac{13\pi}{9})\big)"

"=\sqrt[3]{2}\big(-\cos(\dfrac{4\pi}{9})-i\sin(\dfrac{4\pi}{9})\big)"